Find the kinematic viscosity of an oil having density 980 kg/m^2 when at a certain point in the oil, the shear stress is 0.25 N/m^2 and velocity gradient is 0.3/s .
1
Expert's answer
2020-11-09T19:55:59-0500

By the Newton's Law of Viscosity

"\tau=\mu\dfrac{du}{dy}"

Kinematic viscosity is defined as the ratio between the dynamic viscosity and and density of fluid.


"\nu=\dfrac{\mu}{\rho}"

Sybstitute "\mu=\dfrac{\tau}{\dfrac{du}{dy}}"


"\nu=\dfrac{\tau}{\rho\cdot\dfrac{du}{dy}}"

Given "\rho=980kg/m^3,\tau=0.25N/m^2 , \dfrac{du}{dy}=0.3/s"


"\nu=\dfrac{0.25N/m^2}{980kg/m^3\cdot0.3s^{-1}}\approx0.00085m^2/s=8.5St"

Kinematic viscosity is 8.5 St.